# Mean first passage times of velocity jump processes in higher dimensions

### 31 Mar 2026 v1

@@ Abstract
Velocity jump processes capture a wide range of realistic non-diffusive behaviors [1, 2]. Examples include molecular collisions in biochemistry [3], run-and-tumble motion in bacteria [4, 5, 6, 7, 8], pricing in quantitative finance [9, 10, 11], animal movement [12, 13, 14, 15]. In all cases, the items of interest “run” along a given direction for a random time until a new velocity is selected. The resulting dynamics is intermittent directed motion that can significantly differ from standard diffusion due to the inherent persistence. Velocity jumps may arise from internal mechanisms, external perturbations, or in response to sensory cues. The best known representation of this paradigm are the one-dimensional telegrapher equations that are commonly used to study run-and-tumble behaviors [16, 17, 18, 19]. Non-equilibrium first-passage properties of jump processes are of great significance and may be even more important than steady-state quantities such as moments of the mean displacement or particle distributions [20]. For example, the mean first passage time (MFPT), defined as the average time to first reach a target [21], informs how long one must wait for downstream, often irreversible, events to be initiated, such as immune responses, biochemical transformations, entry into subcellular compartments [22, 23], escape from an enclosed domain [24, 25, 26]. The time to reach a threshold is also used as a measure of resilience in ecology [27] or to quantify the onset of mental health disorders [28]. Understanding how persistence and directionality affect first passage phenomena is critical [29, 30, 31, 32, 33]; however, the problem is notoriously difficult in higher dimensions because velocities can reorient along a continuum of directions, and not just to the left or right as in $d=1$ [34, 35, 36].
@@
Here, $\hat{\theta}=(\cos\theta,\sin\theta)$, $I_{i}(\cdot)$ is the modified Bessel function of the first kind of order $i$, $\hat{\gamma}({\bf{x}})=(\cos\gamma({\bf{x}}),\sin\gamma({\bf{x}}))$ is the preferred direction at $\bf{x}$}, and $\kappa({\bf{x}})\geq 0$ is a unitless concentration parameter, which quantifies the degree of bias towards $\hat{\gamma}({\bf{x}})$. The diffusion tensor and drift are [46] {subequations}
@@
Of particular interest is the narrow capture problem, where a particle seeks a small target in a larger domain [48]. We assume that the particle is initiated at $\rho<r<R_{0}$ and follows the von Mises ($d=2$) or Fisher ($d=3$) distributions with uniform $\kappa$ (and uniform $\alpha,\beta$ as per Table 1) under radial bias. We first take the bulk limit $\ell/\rho\to 0^{+}$, so that for $r\gtrsim\rho+\ell$, $T(r)$ is found by solving Eq. 16 with $z=1$ and $T(\rho)=T^{\prime}(R_{0})=0$. We then take the narrow-capture limit $\rho/R_{0}=\zeta\to 0^{+}$, effectively enlarging the bulk and further suppressing boundary effects. The derived $T(r)$ is used to approximate the global MFPT $\tau_{\rm g}$, defined as spatial average of the MFPT, $\tau_{\rm g}\sim|{\cal{D}}|^{-1}\int_{\cal{D}}T(r)r^{d-1}dr$, leading to the scaling form
@@ Acknowledgements.
We thank T. Chou and K. Painter for fruitful discussions. This work was supported by the ARO (grant W911NF-23-1-0129, MRD), the NSF (grants DMS-2052636, AEL and OAC-2320846, MRD) and the NSERC (Discovery grant RGPIN-2023-04269, TH). Data are available from the authors upon reasonable request .

### 03 Apr 2026 v2

# Mean first passage times of higher-dimensional velocity jump processes

@@ Abstract
Velocity jump processes capture a wide range of realistic non-diffusive behaviors [1, 2]. Examples include molecular collisions in biochemistry [3], run-and-tumble motion in bacteria [5, 6, 7, 8, 4], pricing in quantitative finance [11, 10, 9], animal movement [14, 15, 12, 13]. In all cases, the items of interest “run” along a given direction for a random time until a new velocity is selected. The resulting dynamics is intermittent directed motion that can significantly differ from standard diffusion due to the inherent persistence. Velocity jumps may arise from internal mechanisms, external perturbations, or in response to sensory cues. The best known representation of this paradigm are the one-dimensional telegrapher equations that are commonly used to study run-and-tumble behaviors [16, 18, 19, 17]. Non-equilibrium first-passage properties of jump processes are of great significance and may be even more important than steady-state quantities such as moments of the mean displacement or particle distributions [20]. For example, the mean first passage time (MFPT), defined as the average time to first reach a target [21], informs how long one must wait for downstream, often irreversible, events to be initiated, such as immune responses, biochemical transformations, entry into subcellular compartments [22, 23, 24], escape from an enclosed domain [26, 25, 27]. The time to reach a threshold is also used as a measure of resilience in ecology [28] or to quantify the onset of mental health disorders [29]. Understanding how persistence and directionality affect first passage phenomena is critical [31, 34, 33, 30, 32]; however, the problem is notoriously difficult in higher dimensions because velocities can reorient along a continuum of directions, and not just to the left or right as in $d=1$ [37, 35, 36].
@@
Here, $\hat{\theta}=(\cos\theta,\sin\theta)$, $I_{i}(\cdot)$ is the modified Bessel function of the first kind of order $i$, $\hat{\gamma}({\bf{x}})=(\cos\gamma({\bf{x}}),\sin\gamma({\bf{x}}))$ is the preferred direction at $\bf{x}$}, and $\kappa({\bf{x}})\geq 0$ is a unitless concentration parameter, which quantifies the degree of bias towards $\hat{\gamma}({\bf{x}})$. The diffusion tensor and drift are {subequations}
@@
Of particular interest is the narrow capture problem, where a particle seeks a small target in a larger domain [12, 49]. We assume that the particle is initiated at $\rho<r<R_{0}$ and follows the von Mises ($d=2$) or Fisher ($d=3$) distributions with uniform $\kappa$ (and uniform $\alpha,\beta$ as per Table 1) under radial bias. We first take the bulk limit $\ell/\rho\to 0^{+}$, so that for $r\gtrsim\rho+\ell$, $T(r)$ is found by solving Eq. 16 with $z=1$ and $T(\rho)=T^{\prime}(R_{0})=0$. We then take the narrow-capture limit $\rho/R_{0}=\zeta\to 0^{+}$, effectively enlarging the bulk and further suppressing boundary effects. The derived $T(r)$ is used to approximate the global MFPT $\tau_{\rm g}$, defined as spatial average of the MFPT, $\tau_{\rm g}\sim|{\cal{D}}|^{-1}\int_{\cal{D}}T(r)r^{d-1}dr$, leading to the scaling form
@@ Acknowledgements.
We thank T. Chou and K. Painter for fruitful discussions. This work was supported by the ARO (grant W911NF-23-1-0129, MRD), the NSF (grants DMS-2052636, AEL and OAC-2320846, MRD) and the NSERC (Discovery grant RGPIN-2023-04269, TH). The data that support the findings of this article are openly available [68].